OpenStudy (anonymous):

Help! Consider y = -2x^2+4x-1

OpenStudy (dtan5457):

Solving for x, factoring?

OpenStudy (anonymous):

1.) This equation is given in: (A) Standard form (B) Vertex form (C) Intercept form (D) No particular form 2.) This parabola opens: (A) upward (B) right (C) downward (D) left 3.) This parabola will have (A) a maximum (B) a minimum (C) neither a maximum or a minimum 4.) This axis of symmetry is (A) x=4 (B) x=-2 (C) x=1 (D) x=-1

OpenStudy (anonymous):

My answers: 1.) A 2.) B 3.) B 4.) C

OpenStudy (anonymous):

2 and 3 is incorrect

OpenStudy (dtan5457):

You got 1 and 4 correct

OpenStudy (dtan5457):

consider ax^2+bx+c if a>0, it will open upwards if a<0, it will open downwards downwards parabola=maximum point upwards parabola=minimum point

OpenStudy (anonymous):

So #2 is C, and #3 is A

OpenStudy (dtan5457):

yes

OpenStudy (anonymous):

Thank you!(:

OpenStudy (dtan5457):

welcome